OpenStudy (anonymous):

-3-5(3x+7)=0 ((show work))

OpenStudy (anonymous):

7 times 5 isn't 21?

OpenStudy (anonymous):

the problem should equal positive or negative 38/15

OpenStudy (lxelle):

7 x 5 is 35

OpenStudy (lxelle):

sorry.

OpenStudy (triciaal):

@firelordjohana good catch but the process is correct

OpenStudy (lxelle):

-3-15x-35=0 -15x=38 here you go.

OpenStudy (anonymous):

thank you i got it now!

OpenStudy (triciaal):

distribute, group like terms then simplify